# Hanbury Brown-Twiss interference with massively parallel spectral multiplexing for broadband light

### v2

@@ 1 Introduction
In the following, we describe the results demonstrating the HBT effect of two-photon interference for broadband light, post-selecting photons with the same frequency and timestamps as determined by precise measurement in the spectrometer. We then discuss the relevance of the results to several scenarios of two-photon interference with performance projections based on the achieved performance.
@@ 2.1 LinoSPAD2 detector
The key part of our spectrometer is the LinoSPAD2 camera. In the LinoSPAD2 sensor [8, 49], each pixel is a single-photon sensitive avalanche photodiode (SPAD) with a $26.2\times 26.2$ µm2 size and 25.1% fill factor. The sensor consists of a linear $512\times 1$ pixel array with a median dark count rate of $\sim 125$ cps/pixel, cross-talk of 0.2% [39, 40] and final photon detection efficiency of about 30% at 520 nm and 20% at 640 nm for the sensor equipped with microlenses [8, 7] while operating at room temperature. After calibrations, the temporal resolution was determined to be equal to 40 ps (rms) employing an SPDC source of photon pairs [41]. In this work, we employed 256 pixels (one half of the sensor). LinoSPAD2 has already demonstrated good performance in detecting the HBT photon bunching signatures in our prior experiments with neon and argon spectral lines [32, 21, 39].
@@ 2.2 Spectrometer
The light entering the spectrometer is polarized and filtered via a $(640\pm 5)$ nm bandpass filter before entering a fiber port. A 1-to-2 50:50 fiber-coupled beamsplitter splits the light into two beams, directing it to the two arms of the dual spectrometer. The outputs of the beamsplitter are connected to adjustable collimators that help with focusing and improving the spectral resolution. Each beam is then guided onto a ruled reflective diffraction grating with 1200 grooves/mm after reflection in a silver-coated mirror. The diffracted linear image of the spectrum within each arm of the dual spectrometer is focused through a 200 mm focal length lens onto the sensor.
Every fiber used in this setup, including the fiber-coupled beamsplitter, is single-mode, and everything except the light sources is located inside a light-tight enclosure. The total system efficiency (including the detector) after entering the fiber port was estimated to be about 10%.
@@ 2.3 Light sources and spectral calibrations
The spectrometer was characterized using the neon emission spectrum [38], which has a large number of narrow lines. Figure 3 shows two neon spectra in the dual spectrometer. Two bright neon lines, 638.3 and 640.2 nm, are prominent in both spectrometers. Note that the wavelength is increasing in opposite directions for the two spectrometers due to the orientation of their optical elements. The two neon lines have well-defined wavelengths and can be used to calibrate the correspondence of the wavelength to position in the sensor. The spectral resolution (rms) and pixel scale of the spectrometer was determined to be equal to 0.04 nm and 0.11 nm/pixel, respectively [32].
@@ 2.4 Data processing pipeline
Acquisition of the dataset used for the analysis took nine hours, with a fraction of the detector live time with open shutter of about 10%. On average, the photon rate in a pixel with maximum occupancy was 570 kHz, while the total photon rate from all 256 pixels was approximately 71 MHz, corresponding to a data rate of $\sim 270$ MB/s and resulting in 2 TB of raw data.
In the data analysis, pairs of pixels from the two different spectrometers were selected, and their timestamps were compared calculating delays between detections of pairs of photons. These timestamp differences are then used to build a histogram of coincidences, where a peak — or enhancement in coincidences — may correspond to either the cross-talk or HBT effects [41]. The preprocessed timestamp differences were saved in independent datasets to separate this part of the analysis from its final phase involving the fitting. This resulted in another 2 TB of preprocessed data.
@@ 3.1 HBT peaks
Figure 4: a–d) Time difference distributions in two spectral bins in the dual spectrometer, four combinations in total (pixels 51 and 218, 90 and 179, 50 and 219, 46 and 223), with examples of different contrast values for HBT peaks determined by the fit. e) Sum of all diagonal combinations where the two frequencies are expected to be equal. The peak contrast as determined by a fitting procedure is equal to $(2.0\pm 0.1)$% and sigma 70 ps; f) HBT peak after filtering with $(656.40\pm 0.05)$ nm narrow band filter.
To estimate the interference contrast in an unbiased manner, we developed a simple fitting procedure with a minimal number of free parameters and applied it uniformly to all combinations of pixel pairs. The HBT peak was fitted with a Gaussian function with its position predicted by the offset calibration [40] and constrained to $\pm 200$ ps of its nominal value. The Gauss sigma ($\sigma$) was constrained to the values from 60 to 80 ps, as its possible interval was determined in independent high-statistics measurements. The normalized background was set equal to one. The height of the peak — or contrast ($C$) — was constrained to the values from -10 to 10%.
Typical cases presented in Figure 4 have the peak contrast values from 1.9 to 3.8%. Figure 4 e) shows the sum of time difference distributions where all bin pairs of equal frequencies with expected non-zero signal are combined. The resulting distribution has a prominent HBT peak with $(2.0\pm 0.1)$% contrast and 70 ps sigma. In this case, non-ideal temporal offset calibrations would decrease the contrast and increase the width due to additional smearing, compared to the fits of individual pairs in Figure 4 a–d). We also note that in this case, we acquired enough statistics to observe that the background is not flat but has a slowly changing shape, which was accounted for by a sine term in the fit for this particular case. Figure 4 d) also shows the HBT peak for the LED broadband light after an ultra-narrowband $(656.40\pm 0.05)$ nm filter was used without a spectrometer in a different set of measurements. Note that the width of the filter approximately matches the bin width in the spectrometer, resulting in similar contrasts. The measured contrast, in this case, is $(3.9\pm 0.1)$%.
@@ 3.2 Matrix of contrasts
Finally, Figure 5 shows the full matrix of contrasts for all combinations of pairs of spectral bins in a matrix of $100\times 100$ pixels, so 10,000 combinations in total. We expect HBT peaks to appear only for the cases of equal frequencies for combinations of spectral bins which belong to the diagonal of this matrix as is indeed apparent here. The contrasts were estimated by fitting the time difference distribution as described in detail in the previous subsection. In the case of the off-diagonal elements, where the HBT peaks were not expected, the contrast was close to zero, while for the diagonal elements with HBT peaks, the contrast is about 3 to 4%. We consider this result a remarkable confirmation of expected behavior for the HBT effect massively multiplexed in frequency, and also a sound quality proof of our instrument.
@@ 4 Discussion
Astrometric resolution: The presented results are an important step forward for the stellar intensity interferometry and quantum-assisted schemes of optical interferometers [13, 3] because the spectrometer can be directly employed for spectral binning in the corresponding experiments. In the case of the tested spectrometer, we can combine about 200 pairs of spectral bins where the contrast is non-zero, corresponding to three adjacent diagonal lines. In the considered two-photon interferometer the opening angle between two stars $\Delta\theta$ is extracted from population of HBT peaks measured for different combinations of detectors in two stations of the interferometer [61]. The associated uncertainty of the opening angle can be expressed as in [61], showing dependence on the most important variables:
@@
It is proportional to wavelength $\lambda$ and inversely proportional to baseline (i.e. distance) between interferometer stations $B$, allowing us to gain from longer baselines. The temporal and spectral resolutions of the spectrometer are accounted for by the visibility $V$, while other photon collection and detection parameters, such as the size of the mirrors and the detection efficiency, are accounted for in the detected rate of photons $\bar{n}$. With $T$ being the length of observation period we also note that the overall $T^{-3/2}$ dependence is more favorable than simply the expected $T^{-1/2}$ gain from photon pair statistics, reflecting the advantage of being able to use the measurement of a fringe oscillation rate due to the Earth rotation as proposed in [61] and elaborated in [13].
The contrasts presented in Figure 5 have the same definition as $V$. In the presented study, we directly measured the visibility for each pair of spectral bins and, therefore, can determine the projected uncertainty $\sigma\left[\Delta\theta\right]$ by combining visibility measurements in multiple bins, assuming that all other parameters in Eq. 4.1 are the same for different spectral bins. Employing 70 spectral bins from the main diagonal would allow us to decrease the astrometric uncertainty by a factor of 8.3, while using extra bins from two adjacent diagonals, which are expected to have non-zero visibility, improved this value to 10.0. We consider this as an important step to demonstrate the advantages of frequency multiplexing and the viability of the approach for this application. We also note that the number of spectral bins can be scaled up by a large factor employing larger and multiple sensors, and that it is not unusual to have tens of thousands of spectral channels on modern astronomical spectrographs [48].
@@
where $\hbar$ is Planck’s reduced constant, $\Delta E$, $\Delta f$, and $\Delta t$ are respectively the standard deviations for measurements of energy, frequency, and time.
In our case of 640 nm wavelength and assuming 40 ps & 40 pm rms resolutions [32], $\Delta f\Delta t$ is a factor of 14.7 larger than the Heisenberg-Gabor limit, interpreting the spectral resolution uncertainty as frequency uncertainty. In general, this comparison is a useful benchmark for the indistinguishability of photons if they hit the same spectral bins in the two spectrometers. Experimentally, this means that the contrast of the corresponding HBT peak cannot be larger than $100\%/14.7=6.8\%$. The measured contrast of about 3-4% is indeed limited by the finite spectral and temporal resolutions of the spectrometer.
@@
Given excellent timing resolution of the spectrometer, this architecture is particularly compatible with phase-encoded time-bin qubits [46], in which each qubit is encoded as a photon in a coherent superposition of early and late time bins. Entanglement swapping is then identified by two-photon coincidence events within the same spectral bin [16], which inform the remote parties of the spectral channel hosting the successfully entangled pair for subsequent quantum operations.
@@
Future advancements in fast sensor technologies are expected to significantly enhance temporal resolution, potentially achieving resolutions as fine as 5 ps rms for SPAD devices, as recently demonstrated in [26]. Hybrid solutions with a SPAD sensor attached to a time-stamping chip are also being investigated as a route to large two-dimensional fast imagers [44, 29, 45, 9]. Similarly, SNSPDs are rapidly advancing toward temporal resolutions in the single-digit picosecond regime. However, these detectors currently require cryogenic temperatures, and their scalability to multi-channel and two-dimensional detector arrays has not yet been fully established [10, 37, 54].
@@
Data availability

### v3

@@ 1 Introduction
In the following, we describe the results demonstrating the HBT effect of two-photon interference for broadband light, post-selecting photons with the same frequency and timestamps as determined by precise measurement in the spectrometer. We then discuss the relevance of the results to several scenarios of two-photon interference with projections based on the achieved performance.
@@ 2.1 LinoSPAD2 detector
The key part of our spectrometer is the LinoSPAD2 camera. In the LinoSPAD2 sensor [8, 49], each pixel is a single-photon avalanche photodiode (SPAD) with a $26.2\times 26.2$ µm2 size and 25.1% native fill factor (can be improved with microlenses). The sensor consists of a linear $512\times 1$ pixel array with a median dark count rate of $\sim 125$ cps/pixel, cross-talk of 0.2% [39, 40] and final photon detection efficiency of about 30% at 520 nm and 20% at 640 nm for the sensor equipped with microlenses [8, 7] while operating at room temperature. After calibrations, the temporal resolution was determined to be equal to 40 ps (rms) employing an SPDC source of photon pairs [41]. In this work, we employed 256 pixels (one half of the sensor). LinoSPAD2 has already demonstrated good performance in detecting the HBT photon bunching signatures in our prior experiments with neon and argon spectral lines [32, 21, 39].
@@ 2.2 Spectrometer
The light entering the spectrometer is polarized and filtered via a $(640\pm 5)$ nm bandpass filter before entering a fiber port. A 1-to-2 50:50 fiber-coupled beamsplitter splits the light into two beams, directing it to the two arms of the dual spectrometer. The outputs of the beamsplitter are connected to adjustable collimators that help with focusing and improving the spectral resolution. Each beam is then guided onto a ruled reflective diffraction grating with 1200 grooves/mm after reflection in a silver-coated mirror. The diffracted linear image of the spectrum within each arm of the dual spectrometer is focused through a 200 mm focal length lens onto the sensor. Every fiber used in this setup, including the fiber-coupled beamsplitter, is single-mode, and everything except the light sources is located inside a light-tight enclosure. The total system efficiency (including the detector) after entering the fiber port was estimated to be about 10%.
@@ 2.3 Light sources and spectral calibrations
The spectrometer was characterized using the neon emission spectrum [38], which has a large number of narrow lines. Figure 3 shows two neon spectra in the dual spectrometer. Two bright neon lines, 638.3 and 640.2 nm, are prominent in both spectrometers. Note that the wavelength is increasing in opposite directions for the two spectrometers due to the orientation of their optical elements. The two neon lines have well-defined wavelengths and can be used to calibrate the correspondence of the wavelength to position in the sensor. The resolution of the spectrometer as well as its scale was determined in a separate study [32], where we used the Ar spectrum that we projected onto the sensor of the LinoSPAD2 camera. Fitting with a Gaussian function all of the spectral lines that we measured, we determined the resolution of 0.04 nm (rms), while the scale was estimated using the distances between several known lines and was measured to be equal to 0.11 nm/pixel. Repeating the same procedure with Ne lines, in this work, we measured the same numbers — as expected, since the setup is the same besides the change of the source.
@@ 2.4 Data processing pipeline
Acquisition of the dataset used for the analysis took nine hours, with a fraction of the detector live time with open shutter of about 10%. The time-averaged (not accounting for downtime between acquisition cycles and for data transfer) maximum per-pixel photon count rate was 570 kHz (see Figure 3), while the total photon rate from all 256 pixels was approximately 71 MHz, corresponding to a data rate of $\sim 270$ MB/s and resulting in 2 TB of raw data.
In the data analysis, pairs of pixels from the two different spectrometers were selected, and their timestamps were compared, calculating delays between detections of pairs of photons. These timestamp differences are then used to build a histogram of coincidences, where a peak — or enhancement in coincidences — may correspond to either the cross-talk or HBT effects [41]. The preprocessed timestamp differences were saved in independent datasets to separate this part of the analysis from its final phase involving the fitting. This resulted in another 2 TB of preprocessed data.
@@ 3.1 HBT peaks
Figure 4: a–d) Time difference distributions in two spectral bins in the dual spectrometer, four combinations in total (pixels 51 and 218, 90 and 179, 50 and 219, 46 and 223), with examples of different contrast values for HBT peaks determined by the fit. e) Averaged sum of all diagonal combinations where the two frequencies are expected to be equal. The peak contrast as determined by a fitting procedure is equal to $(2.0\pm 0.1)$% and sigma 70 ps; f) HBT peak after filtering with $(656.40\pm 0.05)$ nm narrow band filter.
To estimate the interference contrast in an unbiased manner, we developed a simple fitting procedure with a minimal number of free parameters and applied it uniformly to all combinations of pixel pairs. The HBT peak was fitted with a Gaussian function with its position predicted by the offset calibration [40] and constrained to $(-5200\pm 200)$ ps of its nominal value. The Gauss sigma ($\sigma$) was constrained to the values from 60 to 80 ps, as its possible interval was determined in independent high-statistics measurements. The normalized background was set equal to one. The height of the peak — or contrast ($C$) — was constrained to the values from -10 to 10%.
Typical cases presented in Figure 4 have the peak contrast values from 1.9 to 3.8%. Figure 4 e) shows the averaged sum of time difference distributions where all bin pairs of equal frequencies with expected non-zero signal are combined. The resulting distribution has a prominent HBT peak with $(2.0\pm 0.1)$% contrast and 70 ps sigma. In this case, non-ideal temporal offset calibrations would decrease the contrast and increase the width due to additional smearing, compared to the fits of individual pairs in Figure 4 a–d). We also note that in this case, we acquired enough statistics to observe that the background is not flat but has a slowly changing shape, which was accounted for by a sine term in the fit for this particular case. Figure 4 f) also shows the HBT peak for the LED broadband light after an ultra-narrowband $(656.40\pm 0.05)$ nm filter was used without a spectrometer in a different set of measurements. Note that the width of the filter approximately matches the bin width in the spectrometer, resulting in similar contrasts. The measured contrast, in this case, is $(3.9\pm 0.1)$%.
@@ 3.2 Matrix of contrasts
Finally, Figure 5 shows the full matrix of contrasts for all combinations of pairs of spectral bins in a matrix of $100\times 100$ pixels, so 10,000 combinations in total. The same fitting procedure that was used for Fig. 4 a)–d) was applied to all pairs of pixels shown here with the same parameter estimation and limitations. We expect HBT peaks to appear only for the cases of equal frequencies for combinations of spectral bins that belong to the diagonal of this matrix, as is indeed apparent here. The contrasts were estimated by fitting the time difference distribution as described in detail in the previous subsection. In the case of the off-diagonal elements, where the HBT peaks were not expected, the contrast was close to zero, while for the diagonal elements with HBT peaks, the contrast is about 2 to 4%. We consider this result a remarkable confirmation of expected behavior for the HBT effect massively multiplexed in frequency, and also a sound quality proof of our instrument.
@@ 4 Discussion
Astrometric resolution: The presented results are an important step forward for the stellar intensity interferometry and quantum-assisted schemes of optical interferometers [13, 3] because the spectrometer can be directly employed for spectral binning in the corresponding experiments, with each bin effectively performing an independent measurement, which can be combined with other measurements to reduce the uncertainty. In the case of the tested spectrometer, we can combine about 200 pairs of spectral bins where the contrast is non-zero, corresponding to three adjacent diagonal lines. In the considered two-photon interferometer, the opening angle between two stars $\Delta\theta$ is extracted from the population of HBT peaks measured for different combinations of detectors in two stations of the interferometer [61]. The associated uncertainty of the opening angle can be expressed as in [61], showing dependence on the most important variables:
@@
It is proportional to wavelength $\lambda$ and inversely proportional to baseline (i.e., distance) $B$ between interferometer stations, allowing us to gain from longer baselines. The temporal and spectral resolutions of the spectrometer are accounted for by the visibility $V$, while other photon collection and detection parameters, such as the size of the mirrors and the detection efficiency, are accounted for in the detected rate of photons $\bar{n}$. With $T$ being the length of observation period we also note that the overall $T^{-3/2}$ dependence is more favorable than simply the expected $T^{-1/2}$ gain from photon pair statistics, reflecting the advantage of being able to use the measurement of a fringe oscillation rate due to the Earth rotation as proposed in [61] and elaborated in [13].
The contrasts presented in Figure 5 have the same definition as $V$. In the presented study, we directly measured the visibility for each pair of spectral bins and, therefore, can determine the improved uncertainty $\sigma\left[\Delta\theta\right]$ by combining visibility measurements in multiple bins, assuming that all other parameters (baselines, photon rates, observing time) in Eq. 4.1 are the same for different spectrometers in both observing stations. Employing 70 spectral bins from the main diagonal would allow us to decrease the astrometric uncertainty by a factor of 8.3, while using extra bins from two adjacent diagonals, which are expected to have non-zero visibility because of the partial spectral overlap of the neighboring spectral bins, improved this value to 10.0. We consider this an important step to demonstrate the advantages of frequency multiplexing and the viability of the approach for this application. We also note that the number of spectral bins can be scaled up by a large factor employing larger and multiple sensors, and that it is not unusual to have tens of thousands of spectral channels on modern astronomical spectrographs [48].
@@
where $\hbar$ is Planck’s reduced constant, $\Delta E$, $\Delta f$, and $\Delta t$ are respectively the standard deviations for measurements of energy, frequency, and time. In our case of 640 nm wavelength and assuming 40 ps & 40 pm rms resolutions [32], $\Delta f\Delta t$ is a factor of 14.7 larger than the Heisenberg-Gabor limit, interpreting the spectral resolution uncertainty as frequency uncertainty. In general, this comparison is a useful benchmark for the indistinguishability of photons if they hit the same spectral bins in the two spectrometers. Experimentally, this means that the contrast of the corresponding HBT peak cannot be larger than $100\%/14.7=6.8\%$, see a complete mathematical treatment of the contrast derivation in Appendix A. The measured contrast of about 3% is indeed limited by the finite spectral and temporal resolutions of the spectrometer, and it can be further diluted by the relative calibrations of the spectrometers and other imperfections.
@@
Given excellent timing resolution of the spectrometer, this architecture is particularly compatible with phase-encoded time-bin qubits [46], in which each qubit is encoded as a photon in a coherent superposition of early and late time bins. Entanglement swapping is then identified by two-photon coincidence events within the same spectral bin [16], which inform the remote parties of the spectral channel hosting the successfully entangled pair for subsequent quantum operations.The main limitation of the presented spectrometer for this application is its low overall system efficiency, estimated at approximately 10% and dominated by the SPAD photon detection efficiency. For two-photon processes, this limitation is particularly severe because the relevant coincidence rates scale with the square of efficiency. Nevertheless, these results represent an important step toward spectral multiplexing for entanglement swapping, demonstrating the feasibility of the approach. As detector technologies continue to improve, higher-efficiency implementations are expected to make this method a competitive and practical solution.
@@
Future advancements in fast sensor technologies are expected to significantly enhance temporal resolution, potentially achieving resolutions as fine as 5 ps rms for SPAD devices, as recently demonstrated in [26]. Hybrid solutions with a SPAD sensor attached to a time-stamping chip are also being investigated as a route to large two-dimensional fast imagers [44, 29, 45, 9]. Similarly, SNSPDs are rapidly advancing toward temporal resolutions in the single-digit picosecond regime. A notable advantage of this type of detector is its superior quantum efficiency compared to SPADs, which is important in many applications. However, these detectors currently require cryogenic temperatures, and their scalability to multi-channel and two-dimensional detector arrays has not yet been fully established [10, 37, 54].
@@
5 Conclusion
We have demonstrated the first observation of the HBT effect for broadband light dispersed in a fast spectrometer across a wide, continuous spectral range of 10 nm, utilizing a large number, of the order of 100, of spectral channels. After precise calibration, our spectrometer achieved exceptional temporal and spectral resolutions of 40 ps and 40 pm, respectively. This enabled the observation of thermal photon bunching with an average contrast of approximately 3%, consistent with theoretical expectations. The achieved temporal and spectral filtering approaches the Heisenberg-Gabor limit, making our spectrometer highly applicable to diverse fields such as quantum communications, quantum metrology, spectroscopy, and quantum-assisted telescopy.
Analyzing the spectrometer’s performance specifically in the context of phase-sensitive intensity interferometry, we found that combining about 200 spectral bin combinations exhibiting non-zero contrast could enhance astrometric precision by a factor of 10. Furthermore, we have outlined the significant potential gains achievable through entanglement swapping in quantum networks, where frequency multiplexing could dramatically enhance the throughput and robustness of entangled photon distribution.
Future advancements in temporal and spectral resolution promise further significant improvements, potentially enabling near-ideal visibility of two-photon interference phenomena. Ultimately, our realization of massively frequency-multiplexed two-photon correlation measurements presents a transformative advancement, with the potential to accelerate the development of quantum sensing and quantum communication technologies by orders of magnitude.
Appendix A Appendix: Spectrally and temporally resolved thermal two-photon correlations
We model the broadband source as a stationary multimode thermal state with
$\langle\hat{a}^{\dagger}(\omega)\hat{a}(\omega^{\prime})\rangle=\overline{n}\,S(\omega)\,\delta(\omega-\omega^{\prime}),\qquad\int_{-\infty}^{\infty}S(\omega)\,d\omega=1,$ (A.1)
where $\overline{n}$ is the total count rate and $S(\omega)$ is the normalized source spectrum, $\hat{a}^{\dagger}(\omega)$ and $\hat{a}$ are standard quantum optical creation and annihilation operators of a photon with frequency $\omega$.
To distinguish the coarse optical filtering performed before the spectrometer from the intrinsic spectral resolution of the detector array, we introduce two different transfer functions:
  • $B(\omega)$ — amplitude transfer function of the external optical filter placed before the detector;
  • $H_{j}(\omega)$ — amplitude spectral response of the $j$-th spectrometer channel (or detector bin).
The effective detected field operator in channel $j$ is then
$\hat{A}_{j}(t)=\int_{-\infty}^{\infty}d\omega\,K_{j}(\omega)\hat{a}(\omega)e^{-i\omega t},\qquad K_{j}(\omega)=B(\omega)H_{j}(\omega).$ (A.2)
The mean count rate in channel $j$ is
$N_{j}=\langle\hat{A}_{j}^{\dagger}(t)\hat{A}_{j}(t)\rangle=\overline{n}\int_{-\infty}^{\infty}S(\omega)|K_{j}(\omega)|^{2}\,d\omega.$ (A.3)
This equation allows one to estimate how many photons are lost due to filtration. For two channels $j$ and $k$, the first-order correlation function is
$G^{(1)}_{jk}(\tau)=\langle\hat{A}_{j}^{\dagger}(t)\hat{A}_{k}(t+\tau)\rangle=\overline{n}\int_{-\infty}^{\infty}S(\omega)\,K_{j}^{*}(\omega)K_{k}(\omega)e^{-i\omega\tau}\,d\omega,$ (A.4)
where $\tau=t_{2}-t_{1}$.
Canonically, thermal sources are usually described via a multimode density operator with a Gaussian-like Glauber-Sudarshan $P$-function, where the fourth-order moment factorizes according to the standard Glauber rule:
$G^{(2)}_{jk}(\tau)=\langle\hat{A}_{j}^{\dagger}(t)\hat{A}_{k}^{\dagger}(t+\tau)\hat{A}_{k}(t+\tau)\hat{A}_{j}(t)\rangle=N_{j}N_{k}+\bigl|G^{(1)}_{jk}(\tau)\bigr|^{2}.$ (A.5)
Therefore, the normalized second-order correlation function is
$g^{(2)}_{jk}(\tau)=\frac{G^{(2)}_{jk}(\tau)}{N_{j}N_{k}}=1+\frac{\bigl|G^{(1)}_{jk}(\tau)\bigr|^{2}}{N_{j}N_{k}}.$ (A.6)
Equation (A.6) immediately explains the main experimental observation: for matched spectral bins ($j=k$), the overlap in Eq. (A.4) is maximal and the HBT bunching peak is observed, whereas for mismatched bins ($j\neq k$), the spectral overlap becomes small and the bunching term is suppressed.
Matched spectral bins.
Below we explore further the diagonal case when $j=k$. Then Eqs. (A.4)–(A.6) reduce to
$g^{(2)}_{jj}(\tau)=1+\left|\frac{\int_{-\infty}^{\infty}S(\omega)|K_{j}(\omega)|^{2}e^{-i\omega\tau}\,d\omega}{\int_{-\infty}^{\infty}S(\omega)|K_{j}(\omega)|^{2}\,d\omega}\right|^{2}.$ (A.7)
Assume now that the source spectrum, the external optical filter, and the detector channel response are Gaussian:
$S(\omega)\propto\exp\!\left[-\frac{(\omega-\omega_{s})^{2}}{2\Delta\omega_{\mathrm{src}}^{2}}\right],$ (A.8)
$|B(\omega)|^{2}\propto\exp\!\left[-\frac{(\omega-\omega_{B})^{2}}{2\delta\omega_{\mathrm{ext}}^{2}}\right],$ (A.9)
$|H_{j}(\omega)|^{2}\propto\exp\!\left[-\frac{(\omega-\omega_{j})^{2}}{2\delta\omega_{\mathrm{det},j}^{2}}\right].$ (A.10)
Then the product
$S(\omega)|B(\omega)|^{2}|H_{j}(\omega)|^{2}$ (A.11)
is again Gaussian, with effective rms width
$\Sigma_{j}^{2}=\left(\frac{1}{\Delta\omega_{\mathrm{src}}^{2}}+\frac{1}{\delta\omega_{\mathrm{ext}}^{2}}+\frac{1}{\delta\omega_{\mathrm{det},j}^{2}}\right)^{-1}.$ (A.12)
Hence Eq. (A.7) becomes
$g^{(2)}_{jj,\mathrm{id}}(\tau)=1+\exp\!\left(-\Sigma_{j}^{2}\tau^{2}\right),$ (A.13)
where $\rm{id}$ indicates an ideal measurement with ideal temporal resolution.
This expression reproduces the expected limits. In particular, if the external optical filter is broad compared to the detector spectral channel width, i.e,
$\delta\omega_{\mathrm{ext}}\gg\delta\omega_{\mathrm{det},j}.$ (A.14)
Then the detector channel dominates the spectral resolution and
$\Sigma_{j}^{2}\approx\left(\frac{1}{\Delta\omega_{\mathrm{src}}^{2}}+\frac{1}{\delta\omega_{\mathrm{det},j}^{2}}\right)^{-1}.$ (A.15)
Finite temporal resolution.
The temporal resolution of the detector array can be modeled by convolving the ideal bunching term with the instrumental response in the delay variable. Let
$R(\tau)=\frac{1}{\sqrt{2\pi}\sigma_{\tau}}\exp\!\left(-\frac{\tau^{2}}{2\sigma_{\tau}^{2}}\right),\qquad\int_{-\infty}^{\infty}R(\tau)\,d\tau=1,$ (A.16)
where $\sigma_{\tau}$ is the rms resolution of the measured time-difference histogram. The measured normalized second-order correlation is then
$g^{(2)}_{jj,\mathrm{M}}(\tau)-1=\int_{-\infty}^{\infty}R(\tau-\tau^{\prime})\left[g^{(2)}_{jj,\mathrm{id}}(\tau^{\prime})-1\right]\,d\tau^{\prime}.$ (A.17)
Substituting Eq. (A.13) into Eq. (A.17) and evaluating the Gaussian convolution gives
$g^{(2)}_{jj,\mathrm{M}}(\tau)=1+\mathcal{C}_{j}\exp\!\left(-\Sigma_{j,\mathrm{M}}^{2}\tau^{2}\right),$ (A.18)
with contrast
$\mathcal{C}_{j}=\frac{1}{\sqrt{1+2\Sigma_{j}^{2}\sigma_{\tau}^{2}}},\qquad\Sigma_{j,\mathrm{M}}=\frac{\Sigma_{j}}{\sqrt{1+2\Sigma_{j}^{2}\sigma_{\tau}^{2}}}.$ (A.19)
If the two spectrometer arms have independent rms timing jitters $\sigma_{t,1}$ and $\sigma_{t,2}$, then
$\sigma_{\tau}^{2}=\sigma_{t,1}^{2}+\sigma_{t,2}^{2}.$ (A.20)
For identical timing resolution in both arms, $\sigma_{t,1}=\sigma_{t,2}=\sigma_{t}$, one obtains
$\sigma_{\tau}=\sqrt{2}\,\sigma_{t},$ (A.21)
and therefore
$\mathcal{C}_{j}=\frac{1}{\sqrt{1+4\Sigma_{j}^{2}\sigma_{t}^{2}}},\qquad\Sigma_{j,\mathrm{M}}=\frac{\Sigma_{j}}{\sqrt{1+4\Sigma_{j}^{2}\sigma_{t}^{2}}}.$ (A.22)
Relation to the Heisenberg–Gabor limit.
In the experimentally relevant regime where the source is broader than a single detector spectral channel and the external optical filter is also broad on that scale, one has
$\Sigma_{j}\approx\delta\omega_{\mathrm{det},j}=2\pi\Delta f,\qquad\mathcal{C}_{j}\approx\frac{1}{\sqrt{1+\left(4\pi\Delta f\,\Delta t\right)^{2}}},$ (A.23)
with asymptotic value for $4\pi\Delta f\,\Delta t\gg 1$:
$\mathcal{C}_{j}\approx\frac{1}{4\pi\Delta f\,\Delta t}.$ (A.24)
Thus, the factor by which the time–frequency resolution product exceeds the Heisenberg–Gabor limit provides a direct upper estimate for the observable HBT contrast in the Gaussian model. For the present experimental setup with $\Delta t=40~\mathrm{ps}$, $\Delta\lambda=40~\mathrm{pm}$ and $\lambda=640$ nm, this factor is equal to 14.7, yielding:
$\mathcal{C}_{j}\approx\frac{100\%}{14.7}\approx 6.8\%.$ (A.25)
Disclosures
Edoardo Charbon is a co-founder of NovoViz. NovoViz has not been involved in this work or in the drafting of this document. The authors declare that they have no conflict of interest.
Code and Data Availability